OpenStudy (anonymous):

The table below shows the distance y, in miles, traveled by a toy car in x minutes: Time (x) (minutes) 10 20 30 40 Distance (y) (miles) 4 8 12 15 Part A: What is the most likely value of the correlation coefficient of the data in the table? Based on the correlation coefficient, describe the relationship between time and distance traveled by the toy car. (4 points)

OpenStudy (anonymous):

Choose the value of correlation coefficient from 1, 0.99, 0.5, 0.02] Part B: What is the value of the slope of the graph of distance versus time between 10 minutes and 30 minutes and what does the slope represent? (3 points) Part C: Does the data in the table represent correlation or causation? Explain your answer. (3 points)

OpenStudy (anonymous):

help please!1111

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

i need help for part A and C

OpenStudy (anonymous):

i got the part b is it the slope=differences between y / differences between x slope=8/20=2/5

OpenStudy (anonymous):

@thu1935

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

@Richard_Feynman

OpenStudy (anonymous):

@Ria23

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

@ganeshie8 please :(

OpenStudy (anonymous):

i need help in just the part A and C

OpenStudy (primeralph):

a. 10 = 4k, k = 2.5 20 = 8k, k = 2.5 again, so 2.5 is a valid corr-coeff. In terms of x, it's 1/2.5 =0.4 c. The data can only represent correlation and not causation. Time cannot cause events, but events can happen at certain times.

OpenStudy (anonymous):

THANKS :)

OpenStudy (ria23):

Beat me to it. D;

OpenStudy (anonymous):

Distance d, time t --> the relation: d/t = 0.4 is a constant , The relationship is: Distance/time = v (speed).